### v2

# SD2: Self-Distilled Sparse Drafters

Abstract
Speculative decoding is a powerful technique for reducing the latency of Large Language Models (LLMs), offering a fault-tolerant framework that enables the use of highly compressed draft models. In this work, we introduce Self-Distilled Sparse Drafters (SD2), a novel methodology that leverages self-data distillation and fine-grained weight sparsity to produce highly efficient and well-aligned draft models. SD2 systematically enhances draft token acceptance rates while significantly reducing Multiply-Accumulate operations (MACs), even in the Universal Assisted Generation (UAG) setting, where draft and target models originate from different model families. On a Llama-3.1-70B target model, SD2 provides a 1.59$\times$ higher Mean Accepted Length (MAL) compared to layer-pruned draft models and reduces MACs by over 43.87% with a 8.36% reduction in MAL compared to a dense draft models. Our 1.5B and 3B unstructured sparse drafters outperform both dense and layer-pruned models in terms of end-to-end latency improvements; highlighting the potential of sparsity-aware fine-tuning and compression strategies to improve LLM inference efficiency while maintaining alignment with target models.
Correspondence to mklasby@ucalgary.$\dagger$ Work completed while on internship at Cerebras
1 Introduction
Large Language Models (LLMs) have proven to have high utility in a wide variety of contexts. However, the causal dependency between preceding and subsequent tokens results in approximately 10$\times$ higher latency for sequence generation compared to processing an equivalent length sequence in parallel (Liu et al., 2023). The high computational cost of LLMs has motivated significant research into methods which improve their efficiency, including: quantization (Gholami et al., 2022; Kurtic et al., 2024), pruning (Ma et al., 2023; Gromov et al., 2025), weight sparsity (Frantar and Alistarh, 2023; Sun et al., 2024b; Yin et al., 2024b), activation sparsity (Mirzadeh et al., 2023), KV-cache compression (Zhang et al., 2024c), distillation (Kim and Rush, 2016; Hsieh et al., 2023), and matrix decomposition (Hu et al., 2021; Liu et al., 2024b). However, these methods typically trade improved efficiency for decreased model quality. (Yin et al., 2024a; Jaiswal et al., 2024).
In contrast, speculative decoding (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023) offers a unique framework to accelerate token generation without sacrificing accuracy. In speculative decoding, a smaller draft model is utilized to auto-regressively generate a sequence of draft tokens which are verified in parallel by a target model. For speculative decoding to be effective, an efficient draft model which is closely aligned with the target model is required. How best to select and/or train a draft model has been the focus of several recent works, see Xia et al. (2024) and Section 4 for more details.
Figure 1: Improvement factor of dense, layer-pruned, and SD2 unstructured Llama and Qwen models drafting for Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct, respectively. SD2 drafters outperform layer-pruned draft models and dense drafters in the 1.5B and 3B model size categories.
(a) UAG Qwen drafting for Llama
(b) MACs analysis
Figure 2: (2(a)): SpecBench MAL for SD2 Qwen-2.5 models drafting for Llama-3.1-70B-Instruct in the UAG setting. These results illustrate the benefits of SD2 for aligning draft models even across different model families. SD2 Qwen drafters achieve a higher MAL than their dense counterparts. (2(b)): MAL vs. MACs for layer-pruned and sparse draft models. Particularly notable are the Qwen-2.5 unstructured sparse drafters which approach iso-MAC performance compared to the dense models.
Fine-grained sparsity, such as unstructured or 2:4 (Mishra et al., 2021) sparsity, for compressing draft models has yet to be examined. Sparse Neural Networks (SNNs) have significantly reduced connectivity between neurons in adjacent layers compared to dense networks. In unstructured sparsity the active parameters are distributed in an irregular, non-uniform manner throughout the weight matrices which can be challenging to accelerate. To address this, 2:4 sparsity was introduced which offers a hardware-friendly structure in which exactly two out of every four contiguous weights are active. However, the additional constraints of 2:4 sparsity result in lower accuracy on downstream tasks compared to unstructured.
Speculative decoding is a uniquely well-suited setting for using sparse draft models since any errors produced during drafting can be gracefully recovered from during the verification step. Other compression strategies such as quantization succeed in reducing memory overhead, but may not reduce latency (Kurtic et al., 2024). Performing operations with parameters quantized to low bit-width data types requires dequantization to a hardware-native data type, typically 8- or 16-bit types for presently available accelerators. As a result, quantization may lead to increased latency in many circumstances, particularly for relatively small draft models which may not be memory-bound. In contrast, SNNs reduce the total amount of Multiply-Accumulate operations (MACs) required per forward pass which, in an ideal setting, would correspond directly with latency improvements. In practice, accessing non-contiguous sparse parameters and storing the associated sparse structure metadata can lead to significant overhead on Graphics Processing Units (GPUs) (Hooker, 2020). Achieving real-world acceleration for fine-grained SNNs requires specialized kernels (Neural Magic, 2021; Schultheis and Babbar, 2023; Lasby et al., 2023; Frantar et al., 2025) and hardware such as Cerebras’ wafer-scale engine (Thangarasa et al., 2024; Hall et al., 2023; Lie, 2023; Agarwalla et al., 2024).
Given the practical limitations of fine-grained sparsity, structured pruning — the removal of layers, neurons, or other substructures — of LLMs to obtain performant draft models is a more hardware-friendly alternative. In particular, pruning entire transformer blocks from the decoder of LLMs provides low latency models which retain their quality under moderate compression ratios (Liu et al., 2023; Gromov et al., 2025; Men et al., 2024; Kim et al., 2024; Sun et al., 2025). A natural extension of these works is to leverage layer-pruning techniques to obtain draft models as examined by  Thangarasa et al. (2025). However, it is not clear a priori if the model quality degradation of layer-pruned models can be overcome by their improved latency for use as draft models in the speculative decoding setting.
In this work, we propose Self-Distilled Sparse Drafters (SD2)— a novel methodology for obtaining efficient, well-aligned draft models — by utilizing self-data distillation fine-tuning (Yang et al., 2024b; Thangarasa et al., 2025) and weight sparsity. Specifically, we make the following contributions:
  • We introduce SD2, a novel methodology for obtaining fine-grained sparse draft models;
  • We demonstrate the superiority of fine-grained sparsity for accelerating speculative decoding and downstream evaluation tasks compared with layer-pruned models;
  • We showcase the effectiveness of self-data distillation fine-tuning for model alignment, even when aligning with a different model family with Universal Assisted Generation (UAG);
  • When paired with optimized sparse representations, we find that the end-to-end acceleration of speculative decoding with fine-grained sparse draft models is comparable to and in some cases exceeds that of dense draft models, particularly with larger draft model sizes.
2 Method
In this section, we provide preliminaries for speculative decoding and introduce the components of SD2, consisting of self-data distillation, one-shot fine-grained pruning, and sparse fine-tuning. We also present our layer pruning method which serves as a baseline to compare to our sparse drafters.
2.1 Speculative decoding
The original motivation for speculative decoding stems from the observation that a wide degree of variance exists in the difficulty of generating tokens. For instance, completions may require copy-pasting portions of the input or including tokens which do not contribute to the semantic content. However, in typical auto-regressive sampling each token requires precisely the same amount of computation to generate regardless of apparent difficulty. In speculative decoding, draft tokens are produced by auto-regressively sampling from a smaller draft model, $M_{d}$, to produce candidate completions which are verified in parallel by the target model, $M_{t}$. The draft-then-verify procedure is repeated for multiple rounds until the generation outputs the end-of-sequence token or another stopping condition occurs – such as reaching a maximum number of generated tokens.
Formally, given an input sequence $\{x_{1},x_{2},...,x_{n}\}$, the draft model calculates probability distributions for each token in the completion conditioned on the input sequence and any preceding output tokens: $p_{n+j+1}=M_{d}(\{x_{1},...,x_{n},...\tilde{x}_{n+j}\})\forall{j}\in\{1,...,k\}$. From these distributions draft tokens are sampled $\tilde{x}_{n+j}\sim p_{n+j}$ to generate a partial completion $\{\tilde{x}_{n+1},...,\tilde{x}_{n+k}\}$ of $k$ draft tokens. In the verification stage, the target model $M_{t}$ computes the output probabilities for each draft token plus one additional token based on its own output distribution: $q_{j+1}=M_{t}(x_{\leq{n}},\tilde{x}_{\leq{j}})\forall{j}\in\{n,...,n+k+1\}$. For each draft token, $\tilde{x_{j}}$, the token is accepted if a verification condition based on the draft and target model probabilities is satisfied. A variety of sampling and verification schemes have been considered in prior work (Stern et al., 2018; Leviathan et al., 2023; Chen et al., 2023). In our experiments, we use greedy sampling with a strict top-1 verification criterion which guarantees that generated text matches the output of the original target model precisely:
$\tilde{x}_{n+j}=\operatorname*{arg\,max}\;p_{j}\quad\forall{j}\in\{1,...,k\},$ (1)
$x_{n+j}=\begin{cases}\tilde{x}_{n+j}&\text{if $\tilde{x}_{n+j}=\operatorname*{arg\,max}\;q_{n+j}$}\\ \operatorname*{arg\,max}\;q_{n+j}&\text{otherwise}.\end{cases}$ (2)
We model the expected improvement factor to the overall latency of speculative decoding versus sampling from the target model directly as follows:
Improvement $\text{Improvement Factor}=\frac{\text{MAL}}{kc+1},$ (3)
where Mean Accepted Length (MAL) is the average number of tokens accepted per round [^1] MAL includes accepted draft tokens and one additional token from the target model per round., $k$ is the number of draft tokens speculated per round, and $c$ is the cost factor representing the ratio of draft versus target model efficiency. For our practical improvement factor calculations, $c$ represents the wall clock latency ratio, i.e., $c=\frac{\text{timeit}(M_{d}(x))}{\text{timeit}(M_{t}(x))}$. For our theoretical improvement factor calculations, we substitute MACs in lieu of latencies.
Algorithm 1 Self-data distillation for speculative decoding
1:  Input: Pretrained target model $M_{t}$, fine-tuning dataset $D_{f}$ with prompts $\mathcal{X}$, context $\mathcal{C}$, and labels $\mathcal{Y}$
2:  Initialize $D_{self}=\emptyset$
3:  for $\mathbf{X_{i}},\mathbf{Y_{i}},\mathbf{C_{i}}\in D_{f}$ do
4:     $\mathbf{X^{\prime}_{i}}\leftarrow\mathbin{\|}\mathbf{C_{i}}\mathbin{\|}\mathbf{X_{i}}\mathbin{\|}\mathbf{Y_{i}}$Combine into new prompt}
5:     $\mathbf{\tilde{Y_{i}}}\sim M_{t}(\mathbf{X^{\prime}_{i}})$Generate new label}
6:     $D_{self}\leftarrow(\mathbf{X_{i}},\mathbf{\tilde{Y_{i}}})$Accept $\mathbf{\tilde{Y_{i}}}$ w/o verification}
7:  end for
8:  Output: $D_{self}$
2.2 Self-data distillation
Self-data distillation consists of generating a dataset, $D_{self}$, whose outputs are generated by a model of interest. We curate fine-tuning datasets following  Yang et al. (2024b); Thangarasa et al. (2025). The distilled labels are generated by the target model, $M_{t}$ based on the input sequences, ground truth labels, and task-specific context of one or more supervised fine-tuning datasets $D_{f}$. Specifically, given a task specific context $\mathbf{C^{t}}$, original input sequence $\mathbf{X^{t}}$, and original ground truth label $\mathbf{Y^{t}}$. We combine these components into a new input sequence $\mathbf{X^{\prime}}$
$\mathbf{\tilde{Y}}=M_{t}(\mathbf{X^{\prime}})\quad\text{where}\;\mathbf{X}^{\prime}=\mathbf{C^{t}}\mathbin{\|}\mathbf{X^{t}}\mathbin{\|}\mathbf{Y^{t}}.$ (4)
In the original formulation of self-data distillation, Yang et al. (2024b) only extract the distilled label $\mathbf{\tilde{Y}}$ if and only if it aligns with the original ground truth label, ${\mathbf{Y^{t}}}$. For instance, only accepting the distilled label if it is mathematically equivalent to the ground truth label. However, for tasks that do not yield a closed-form solution, it can be challenging to assess the validity of the distilled label.
The speculative decoding setting gracefully eliminates the need to consider the distilled label verification process. Since our fundamental goal is aligning the draft with the target model, the correctness of the distilled label is irrelevant. We accept the distilled labels even if they contain easily identifiable errors; the distilled output, correct or otherwise, is already aligned with the target model output distribution. See Algorithm 1 for a summary of the self-data distillation process.
2.3 Fine-grained pruning
To obtain our fine-grained sparse draft models, $M_{d}$, we use SparseGPT (Frantar and Alistarh, 2023). However, we emphasize that our method is compatible with any pruning algorithm. In addition to the original SparseGPT hyperparameter settings, we experiment with non-uniform layer-wise sparsity distributions such as Outlier Weighted Layer-wise sparsity (OWL) (Yin et al., 2024b) and a novel distribution inspired by the angular cosine distance measure from  Gromov et al. (2025). While our experiments found that non-uniform layer-wise sparsity improves perplexity, little benefit was observed on downstream evaluations or speculative decoding. See Appendix D for a more thorough discussion of our findings. Ultimately, we use a uniform layer-wise sparsity distribution for all results presented below.
2.4 Sparse fine-tuning
In typical supervised fine-tuning settings, a dense LLM $M$ with parameters $\theta$ is fine-tuned using a supervised dataset $D_{f}$ containing input sequences $\mathcal{X}$ and ground truth output sequences $\mathcal{Y}$. For each token, $y_{j}$, in the ground truth output sequences, $\mathbf{Y_{i}}\in\mathcal{Y}$, the model outputs a probability distribution over its vocabulary conditioned by the input sequences, $\mathbf{X_{i}}\in\mathcal{X}$, any output tokens preceding the current token position, and the model parameters: $Q({y}_{j}|\mathbf{X_{i}},\{y_{1},...,y_{j-1}\},\theta)$ where $Q(\cdot)$ represents the language modelling head (LM head) logits normalized with the softmax function. The total loss per mini-batch is the average negative log-likelihood across all sequences and tokens:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{S_{i}}\sum_{j=1}^{S_{i}-1}P(y_{j})\log(Q({y}_{j}|\mathbf{X_{i}},y_{\leq j-1},\theta)),$ (5)
where $P(y_{j})$ is the ground truth distribution, $N$ is the number of samples in the batch, and $S_{i}$ is the output sequence length. The model is fine-tuned by minimizing this loss using stochastic gradient descent. In the sparse fine-tuning setting, we simply replace $\theta$ with $\theta_{s}\subset\theta$, the set of non-zero parameters. In our approach, the sparse topology is fixed after one-shot pruning. Pruned parameters remain fixed at zero throughout fine-tuning.
For our experiments, we fine-tune our sparse models with a binary mask to initialize pruned parameters to zero and set their gradients to zero during backpropagation via backwards hook. The backwards hook ensures that gradients of pruned parameters do not contribute to the partial derivates of active parameters nor optimizer buffer states. See Algorithm 2 for more details.
Algorithm 2 Sparse fine-tuning
  Input: Pruned draft model $M_{d}^{\prime}$ with trainable parameters $\theta$, pruned parameters $\theta_{p}$, self-data distilled dataset $D_{self}=\{(\mathbf{X}_{i},\mathbf{\tilde{Y}}_{i})\}_{i=1}^{T*N}$ with distilled outputs sequences of length $S_{i}$, optimizer $\mathcal{O}$ (e.g., AdamW), learning rate $\alpha$, number of iterations $T$, and batch size $N$.
  Define GradientHook($\theta,\theta_{p},\nabla_{\theta}\mathcal{L}_{t}$):
[…]
     for $p_{i},\frac{\partial\mathcal{L}_{t}}{\partial p_{i}}\in\{\theta,\nabla_{\theta}\mathcal{L}_{t}\}$ do
        if $p_{i}\in\theta_{p}$ then
           $\frac{\partial\mathcal{L}_{t}}{\partial p_{i}}\leftarrow 0$
        end if
     end for
  end if
  return $\nabla_{\theta_{s}}\mathcal{L}_{t}$
  $\theta_{s}\leftarrow\theta\notin\theta_{p}$Set of active parameters}
[…]
  for $t=1$ to $T$ do
     $(\mathbf{X}_{i},\mathbf{\tilde{Y}}_{i})\sim D_{self}$Sample mini-batch}
     $\mathcal{L}_{t}\leftarrow 0$Initialize mini-batch loss}
     for $n=1$ to $N$ do
       ${L}_{n}\leftarrow 0$Initialize sequence loss}
       for $j=1$ to $S_{i}-1$ do
          $L_{n}\mathrel{{+}{=}}P(\tilde{y}_{j})\log Q(\tilde{y}_{j}|\mathbf{X_{i}},\tilde{y}_{\leq j-1},\theta_{s})$
       end for
       $\mathcal{L}_{t}\mathrel{{+}{=}}(L_{n}/(S_{i}-1))/N$
     end for
[…]
     $\theta_{s}\leftarrow\mathcal{O}(\theta_{s},\nabla_{\theta_{s}}\mathcal{L}_{t},\alpha)$
  end for
  Output: $M_{d}^{\prime}$
2.5 Layer pruning
We obtain our layer-pruned baselines as follows. Consider a draft model $M_{d}$ with $N$ decoder blocks. Each decoder block consists of a multi-headed self-attention module followed by a feed-forward module. The decoder blocks are stacked sequentially such that the output of the preceding block is the input to the following block. The final block’s output is typically used for downstream tasks such as language modelling or sequence classification. The goal of layer pruning is to identify and prune $n$ sequential decoder blocks such that the resultant quality degradation on relevant downstream tasks is minimized.
Various approaches have been proposed for identifying the most important blocks in a transformer (Samragh et al., 2023; Men et al., 2024). Following Thangarasa et al. (2025), we elect to use the angular cosine distance measure as proposed by  Gromov et al. (2025). This metric uses the angular cosine distance between the input and output of a group of $n$ sequential decoder blocks to define block importance. The group of blocks with the highest similarity between their input and output are considered to be more redundant than other group candidates and therefore can be pruned with the smallest impact on the overall model output. Formally, we define the angular cosine distance measure as follows:
$d(x_{D_{t}}^{i}\text{,}\;x_{D_{t}}^{i+n})=\frac{1}{\pi}\text{arccos}\left(\frac{x_{D_{t}}^{i}\cdot x_{D_{t}}^{i+n}}{\left\|x_{D_{t}}^{i}\right\|\left\|x_{D_{t}}^{i+n}\right\|}\right),$ (6)
[…]
$i^{*}(n)=\operatorname*{arg\,min}_{i}\;d(x_{D_{t}}^{i}\text{,}\;x_{D_{t}}^{i+n}),$ (7)
where $i^{*}(n)$ is the starting block index for the block group of length $n$ with the minimal angular cosine distance for a given calibration dataset subset. Once identified, decoder blocks $i^{*}$ to $i^{*}+(n-1)$ are pruned and the outputs of layer $i^{*}$ connect as input to block $i^{*}+n$ to obtain the pruned draft model $M_{d}^{\prime}$.
(a) Llama-3.2
(b) Qwen-2.5
Figure 3: Average accuracy on OpenLLM Leaderboard V1 benchmarks for dense, layer-pruned, and sparse Llama-3.2 (Figure 3(a)) and Qwen-2.5 (Figure 3(b)) draft models. Lighter and darker shades show results after one-shot pruning and SD2 fine-tuning, respectively. Self-data distillation with sparse fine-tuning enables recovery of accuracy post-pruning. Notably, Qwen-2.5-3B-Instruct at 50% unstructured sparsity exceeds the accuracy of the original dense model after preparation with SD2
3 Results
To empirically evaluate SD2 and our primary hypothesis that sparse draft models can outperform dense or layer-pruned models, we evaluate a variety of draft model candidates across the Llama-3.2 (Llama Team, AI @ Meta, 2024) and Qwen-2.5 (Yang et al., 2025) model families. For Llama, we evaluate drafting with Llama-3.2 1B-Instruct and 3B-Instruct for Llama-3.1-70B-Instruct. For Qwen-2.5, we evaluate 0.5B-Instruct, 1.5B-Instruct, 3B-Instruct drafting for Qwen-2.5-72B-Instruct. We also consider the UAG setting in which the Qwen draft models are aligned to and evaluated with a Llama-3.1-70B-Instruct target model. We prune and fine-tune fine-grained sparse draft models with both unstructured and 2:4 sparsity using a uniform layer-wise sparsity distribution and prune all decoder blocks, excluding the embedding and LM head. The sparsity levels reported in our results refer to the overall sparsity of the decoder. As baselines, we compare these candidates with 50% layer-pruned [^2] Specifically, 50% of the decoder blocks are pruned. and dense models. See Appendix C for hyperparameters and implementation details.
3.1 Evaluation
We evaluate the draft models for downstream task accuracy on the OpenLLM Leaderboard V1 benchmarks (HuggingFace, 2024) and SpecBench (Xia et al., 2024).
3.1.1 OpenLLM leaderboard V1 benchmarks
We evaluate our models using the default multi-shot settings on the OpenLLM Leaderboard V1 tasks using the EleutherAI evaluation harness (Gao et al., 2023). These tasks include: 25-shot ARC-C (Clark et al., 2018), 5-shot strict exact match GSM8k (Cobbe et al., 2021), 10-shot HellaSwag (Zellers et al., 2019), 5-shot MMLU (Hendrycks et al., 2021), 5-shot Winogrande (Sakaguchi et al., 2021), and 0-shot multi-true (MC2) TruthfulQA (Lin et al., 2022). We report byte-length normalized accuracies for ARC-C and HellaSwag [^3] Reported as the acc_norm field in the EleutherAI evaluation harness outputs. See  Gao (2021) for details..
See Figure 3 for the mean accuracy for models pruned and fine-tuned using the SD2 methodology. We find that one-shot pruned models suffer high degradation on the tasks evaluated; however, after fine-tuning the fine-grained sparse models recover much of their accuracy, particularly the 50% unstructured and 2:4 sparse models. For the Qwen-2.5 1.5 and 3B models, we find that the unstructured 50% models approach or exceed the accuracy of the dense baseline. Notably, the 50% layer-pruned models suffer high degradation, even after fine-tuning. See Appendix G for more detailed results.
(c) Llama-3.2-1B-Instruct
(d) Llama-3.2-3B-Instruct
(e) Qwen-2.5-0.5B-Instruct
(f) Qwen-2.5-1.5B-Instruct
(g) Qwen-2.5-3B-Instruct
Figure 4: MAL for Llama-3.2 and Qwen-2.5 model families on SpecBench (Xia et al., 2024) for layer-pruned, unstructured sparse, and 2:4 sparse draft models including results after one-shot pruning and SD2 fine-tuning. Dense model baseline are depicted in horizontal grey lines on each plot. Top row: Llama-3.2 draft models speculating for a Llama-3.1-70B-Instruct target model. Bottom row: Qwen-2.5 draft models speculating for Qwen-2.5-72B-Instruct. Across both model families, we observe that the fine-grained sparse draft models achieve significantly higher MALs than the layer-pruned models.
3.1.2 Speculative decoding
SpecBench Xia et al. (2024) assesses draft token acceptance rate across a variety of tasks including translation, summarization, question answering, mathematical reasoning, multi-turn conversation, and retrieval augmented generation (RAG).
In Figure 4, we report the MAL on SpecBench (Xia et al., 2024) for a variety of models and sparsities. We find that the SD2 50% unstructured drafters achieve the highest MAL of the sparse models, but do not exceed the dense baselines. For Qwen-2.5, the SD2 50% unstructured drafter achieve a MAL within 0.2 of the dense models across all model sizes investigated, despite lower overall scores compared to the Llama model family. We speculate that the overall higher MAL for the Llama-3.2 draft models are the result of improved alignment with the target model stemming from the pruning and distillation process used during pretraining (Llama Team, AI @ Meta, 2024).
The importance of the draft model alignment is clear; our self-data distillation and fine-tuning process enables much higher MAL than the one-shot pruned models. Even more strikingly, the SD2 effectively aligns the draft model in the UAG setting across model families, as can be seen in Figure 2(a). Remarkably, the sparse Qwen drafters outperform their dense counterparts under this setting, further highlighting the importance of draft model alignment and the benefits of SD2.
3.2 MACs analysis
The prior section demonstrate that SD2 drafters have comparable quality to dense models, it remains unclear whether their efficiency gains yield offer practical benefits. Unstructured sparsity is not well suited for GPUs and tangible latency improvements require specialized kernels. An analysis based on MACs provides a theoretical limit of the benefits one can expect from unstructured sparsity.
Figure 5: Theoretical improvement factor for SD2 Llama-3.2 models drafting for a Llama-3.1-70B-Instruct based on MAL and MAC s. Our SD2 Llama-3B draft model offers the highest improvement factor compared to dense or layer-pruned models.
In Figure 2(b), we plot the MAL of our models trained with SD2 versus their corresponding MACs. The unstructured sparse Qwen-2.5 drafters nearly achieve iso-MAC performance improvements compared to their dense counterparts. However, additional MACs associated with the LM head prevent the unstructured sparse models from directly outperforming the dense drafters in terms of MAL vs. MACs. In Figure 5 we demonstrate that our best performing SD2 draft models provide the highest theoretical improvement factor of all draft model candidates examined.
3.3 Compressed sparse drafters
As noted in Equation 3, a crucial property to consider when evaluating speculative decoding draft models is $c$, the ratio of the draft to target model latency. To realize practical benefits of SD2 on commodity hardware available today, we require efficient sparse representations and kernels which offer lower latency than the dense baselines. To this end, we benchmark our SD2 unstructured draft models using nm-vLLM (Neural Magic, 2024; Kwon et al., 2023) and compare them with dense and layer-pruned draft models in Figure 1. SD2 drafters outperform layer-pruned models and even dense drafters in the 3B model size category. While the smallest dense model in each model family achieves a higher overall improvement factor, these results highlight the potential for sparsity to enable the use of larger draft models. As foundation model sizes continue to grow, increasing the capacity of draft models may be necessary to maintain sufficient MAL and support novel use-cases such as scaling test-time compute. See Appendix E for details on our benchmarking methodology.
4 Related work
Speculative decoding with greedy verification was initially proposed by Stern et al. (2018). Speculative sampling was introduced concurrently by Leviathan et al. (2023) and Chen et al. (2023). Since these initial investigations, several works have proposed modifications and refinements to the original framework. In Medusa (Cai et al., 2024), the authors reformulated used a single model with multiple lm-head layers, each of which predicts multiple draft token candidates across each draft token position, introducing the first instance of draft token-trees. Specifically related to our work, Medusa also used self-distillation to obtain a fine-tuning dataset. Sadhukhan et al. (2024) examined speculative decoding for long context lengths, finding that constant KV-cache sizes lead to improved acceleration. EAGLE (Li et al., 2024c, b) proposed an efficient draft model consisting of a single layer auto-regression head trained from scratch and the frozen embedding and lm-head layers from the target model. Token trees were further investigated by Miao et al. (2024) in which a parallel decoding algorithm was also proposed. Sun et al. (2024a) introduced a hierarchical framework in which efficient KV-cache implementations are used for drafting. Speculative decoding was combined with early-exit mechanisms for drafting by Zhang et al. (2024a) and Liu et al. (2024a). We note that optimizations such as draft token-trees and KV-cache compression are orthogonal to SD2 and future work could consider their integration with our method.
The surprising result that a significant fraction of decoder blocks can be pruned from LLMs without incurring excessive quality reduction was examined in detail in several works (Jha et al., 2024; Sun et al., 2025; Yang et al., 2024a; Men et al., 2024). In particular, these works found that the “middle” decoder blocks are the most amenable to pruning. However, abstract reasoning tasks were noted to be disproportionality affected by layer-pruning. We specifically highlight Gromov et al. (2025) which introduced the angular cosine distance metric that is used for our layer-pruned baselines.
Fine-grained sparsity for LLMs has received significant attention from the research community. Efficient one-shot pruning methods such as SparseGPT (Frantar and Alistarh, 2023), Wanda (Sun et al., 2024b), and Plug-and-Play (Zhang et al., 2024b) proposed re-framing the pruning procedure as a layer-wise reconstruction of the original model weights, using the magnitudes and/or activations to determine the saliency of individual weights. Fang et al. (2024) proposed freezing the model weights and training the mask alone to efficiently obtain 2:4 sparse LLMs. OWL (Yin et al., 2024b) and AlphaPruning (Lu et al., 2024) proposed non-uniform layer-wise sparsity distributions which outperform uniform distributions in terms of perplexity and downstream task accuracy. Despite these efforts, sparse LLMs still fall short of their dense counterparts in many respects (Jaiswal et al., 2024; Thangarasa et al., 2023; Yin et al., 2024a).
5 Conclusion
We introduce the SD2 methodology for obtaining fine-grained sparse draft models. SD2 consists of self-data distillation, one-shot pruning, and sparse fine-tuning. We find that self-data distillation effectively aligns draft models, even if their target is from a different model family such as in the UAG setting. SD2 shows significant theoretical benefits based MACs. When paired with efficient sparse kernels, SD2 outperforms layer-pruned models at all model sizes and compression ratios evaluated. For the 3B draft models, SD2 improves end-to-end speculative decoding latency compared to dense draft models, revealing the potential for using larger drafters. See Appendix B for a discussion of limitations and future work. We hope this work will inspire further efforts to develop low-latency sparse kernels and encourage software/hardware co-design for efficient LLM inference.
6 Acknowledgements
We would like to acknowledge the helpful feedback of Ganesh Venkatesh during our initial explorations that led to this work. ML and YI gratefully acknowledge the support of Alberta Innovates (ALLRP-577350-22, ALLRP-222301502), the Natural Sciences and Engineering Research Council of Canada (NSERC) (RGPIN-2022-03120, DGECR-2022-00358), and Defence Research and Development Canada (DGDND-2022-03120). YI is supported by a Schulich Research Chair.
Appendix A Impact Statement
Training and serving LLMs consume large amounts of energy, even for efficient implementations that leverage techniques discussed in this paper. Speculative decoding requires more VRAM than auto-regressive sampling, necessitating the use of additional hardware devices. There are many potential negative societal consequences of generative AI, we hope that by improving the efficiency of these models we will help distribute the benefits of such systems more equitably to the broader public.
Appendix B Limitations and future work
The effectiveness of SD2 in improving inference efficiency depends on software and hardware optimized for sparsity. The nm-vLLM unstructured sparse kernel we use to benchmark is optimized for Ampere-series NVIDIA GPUs, and evaluating its performance on alternative hardware, such as TPUs or custom accelerators, is crucial for broader adoption. While SD2 requires significant compute for fine-tuning and self-data distillation, this investment could further enhance the quality of dense models, particularly in speculative decoding. Several promising research directions emerge from this work. One is integrating quantization-aware training (Chen et al., 2025), draft-token trees, and compressed KV-cache implementations (Luohe et al., 2024) with SD2 to improve memory and compute efficiency. Another is benchmarking SD2 on hardware optimized for unstructured sparsity, such as Cerebras’ wafer-scale engine. Incorporating advanced fine-tuning strategies like speculative knowledge distillation (Xu et al., 2024a) or square-head distillation (Kurtic et al., 2023) could further improve sparse draft model quality. Finally, combining structured pruning with fine-grained sparsity and quantization presents an opportunity to reduce inference-time latency while maintaining accuracy, making sparsity-aware methods like SD2 more widely applicable across diverse hardware architectures.
Appendix C Hyperparameter settings and implementation details
The core components of SD2 consist of: one-shot pruning, self-data distillation, and sparse fine-tuning.
C.1 One-shot pruning
For one-shot pruning, we use the default SparseGPT hyperparameters. Explicitly, we use 0.01, 128, and 16 bits for the Hessian damping, block size, and model precision, respectively. For calibration, we randomly select 128 samples from a subset [^4] https://huggingface.co/datasets/robbiegwaldd/dclm-micro of the DCLM dataset (Li et al., 2024a) with a sequence length of 2048 tokens. We selected the DCLM calibration dataset based on the results of Ji et al. (2025).
We prune our draft model candidates to 0.5, 0.66, and 0.75 unstructured sparsity in addition to 2:4 sparsity. For our experiments with non-uniform layer-wise sparsity distributions, we explore both OWL and our proposed angular distance layer-wise distribution, the results of which are presented in Appendix D.
However, despite apparent advantages in terms of perplexity when measured on WikiText-V2 (Merity et al., 2017), we find that the non-uniform layer-wise sparsity distributions considered offer little benefit to downstream evaluation tasks or MAL during speculative decoding. Furthermore, since non-uniform layer-wise sparsity distribution pose a challenge for inference systems that leverage pipelining with continuous batching, we opt to use the more straightforward uniform layer-wise sparsity distribution for all results included in Section 3.
C.2 Self-data distillation
Our self-data distillation follows the original implementation [^5] https://github.com/sail-sg/sdft of  Yang et al. (2024b). We generate the distilled labels using 16 bit float precision for the model weights and a maximum generation length of 4096 tokens. We sample from the target model with top-p sampling with $p=1.0$ and a temperature of 0.9. Using Llama-3.1-70B, we use self-data distillation to produce aligned fine-tuning datasets for Opus-100 translation (Tiedemann and Thottingal, 2020) [^6] https://huggingface.co/datasets/Helsinki-NLP/opus-100 and MathInstruct-V2 (Toshniwal et al., 2025) datasets. Using Qwen-2.5-72B-Instruct, we distil MathInstruct-V2.
We create the distillation inputs by combining the original input, original output, and context with the default chat templates. For MathInstruct-V2 we applied the following template:
1_register_template(
2 name="llama3_orig_math_distill",
3 format_user=StringFormatter(
4 slots=[
5 (
[…]
8 )
9 ]
10 ),
[…]
12 format_observation=StringFormatter(
13 slots=[
14 (
[…]
17 )
18 ]
19 ),
20 format_prefix=EmptyFormatter(slots=[{"bos_token"}]),
21 default_system=(
22 "You are a math expert tasked with generating high-quality responses. You will be provided with a math question and a reference answer. Your goal is to rewrite the reference answer in a clear and accurate manner, ensuring it thoroughly addresses the question. Maintain mathematical rigor while improving clarity where necessary."
23 ),
24 stop_words=["<|eot_id|>"],
25 replace_eos=True,
26 replace_jinja_template=False,
27)
For Opus, we applied the following template:
1_register_template(
2 name="llama3_multilingual_distill",
3 format_user=StringFormatter(
4 slots=[
5 (
[…]
8 )
9 ]
10 ),
11 format_system=StringFormatter(
[…]
13 ),
14 default_system=(
15 "You are a language translation expert tasked with generating high-quality translations. You will be provided with a sentence or passage in the source language and a reference translation. Your goal is to rewrite the reference translation to ensure it is both accurate and fluent, preserving the original meaning while improving clarity and readability. Ensure cultural nuances and context are respected during the translation process."
16 ),
17 stop_words=["<|eot_id|>"],
18 replace_eos=True,
19)
In addition to these datasets, we leverage publicly available datasets produced using the self-synthesis method Magpie (Xu et al., 2024b) for both Llama-3.1-70B-Instruct [^7] https://huggingface.co/datasets/Magpie-Align/Magpie-Llama-3.1-Pro-MT-300K-Filtered and Qwen-2.5-72B-Instruct [^8] https://huggingface.co/datasets/Magpie-Align/Magpie-Qwen2.5-Pro-300K-Filtered target models. As the Magpie datasets are generated by the target models, further alignment with self-data distillation is redundant.
C.3 Sparse fine-tuning
For sparse fine-tuning, we extend Llama-Factory Zheng et al. (2024) to incorporate our sparse fine-tuning method. We fine-tune our models using a maximum sequence length of 8,192 tokens and truncate any tokens which exceed this limit. For Llama models, we randomly interleave samples from all three fine-tuning datasets with sampling probabilities 75, 12.5, and 12.5% from Magpie, distilled Opus translation, and distilled MathInstruct-V2, respectively. For Qwen, we randomly interleave samples from Magpie and distilled MathInstruct-V2 with probabilities 80 and 20%, respectively.
We optimize our models using the default AdamW optimizer (Loshchilov and Hutter, 2018) with 1.0e-08, 0.9, and 0.999 for the $\varepsilon$, $\beta_{1}$, and $\beta_{2}$ hyperparameters, respectively. For most of our models, we train them with 16,000 optimizer steps with a batch size of 8 for a total of 128,000 samples. The one exception to the above are the 75% sparse models, which we find benefit from an extended training duration of 32,000 optimizer steps with a batch size of 8 for a total of 256,000 samples.
We use a grid search optimize the learning rate for all Llama draft model and sparsity combinations. In general, we find that smaller and more sparse models required higher learning rates. Rather than performing a second grid search, we simply re-use the optimal learning rates found for Llama models when fine-tuning Qwen models, using the Llama-3.2-1B-Instruct learning rates for both 0.5B and 1.5B Qwen models. See Table 1 for the various learning rates used based on the model size and sparsity type. For learning rate schedule, we use a linear schedule with a linear warm-up using 5% of the total steps. We also experimented with a cosine schedule but found no significant difference in the validation loss.
Table 1: Learning rates used for sparse fine-tuning of the various draft model candidates.
Model Sparsity LR
Llama-3.2-1B-Instruct 0.5 1.25e-05
2:4 1.25e-05
50% layer-pruned 1.25e-05
0.66 2.0e-05
0.75 5.0e-05
Llama-3.2-3B-Instruct 0.5 7.5e-06
2:4 7.5e-06
50% layer-pruned 7.5e-06
0.66 1.25e-05
0.75 2.0e-05
Qwen-2.5-0.5B-Instruct 0.5 1.25e-05
2:4 1.25e-05
50% layer-pruned 1.25e-05
Qwen-2.5-1.5B-Instruct 0.5 1.25e-05
2:4 1.25e-05
50% layer-pruned 1.25e-05
Qwen-2.5-3B-Instruct 0.5 7.5e-06
2:4 7.5e-06
50% layer-pruned 7.5e-06
C.4 Computational Resources
We use an internal cluster to run our experiments on a node with four Nvidia H100-80GB-HBM3 GPUs, 192 Intel Xeon Platinum 8468 CPUs, and 2TB of RAM. Pruning was completed using a single GPU in less than 30 minutes per model. Our fine-tuning datasets were generated in approximately 6-8 hours per dataset. For sparse fine-tuning, we use a single GPU with the following approximate runtimes: 10 and 21 hours for 1B and 3B, respectively, for the 50% and 66% sparse models. The 75% sparse model fine-tuning took approximately 20 and 42 hours for 1B and 3B, respectively, due to the increased number of optimization steps.
Appendix D Non-uniform layer-wise sparsity distributions
This section highlights our analysis of non-uniform layer-wise sparsity distributions such as OWL (Yin et al., 2024b) and a novel distribution inspired by the angular cosine distance measure from  Gromov et al. (2025). Ultimately, we found that the non-uniform layer-wise sparsity distributions examined resulted in improved perplexity on WikiText-V2 and a modest improvement on the OpenLLM Leaderboard V1 suite of benchmarks; however, we did not find a statistically significant improvement to MAL during speculative decoding.
OWL allocates sparsity to layers proportional to their outlier ratio. The outlier ratio was inspired by the observation that the activations of LLMs often contain large outliers (Dettmers et al., 2022). Specifically, for a weight matrix $\mathbf{W}$ of shape ($C_{out}$, $C_{in}$) and input $X$, OWL defines the outlier score of $\mathbf{W}_{i,j}$ as $A_{i,j}=\left\|X_{j}\right\|\cdot\left|\mathbf{W}_{i,j}\right|$ across all tokens in a calibration dataset. The sparsity ratio is defined as $S^{l}\propto 1-D^{l}\quad S^{l}\in\{S-\lambda,S+\lambda\}$ where $S\in[0,1]$ is the uniform sparsity target, $\lambda$ is a hyperparameter which constraints the maximum and minimum layer sparsities, and $D^{l}$ is the outlier distribution for a single layer calculated as
$D^{l}=\frac{\sum_{i=1}^{C_{out}}\sum_{j=1}^{C_{in}}\mathbf{1}_{\zeta}(A^{l}_{i,j})}{C_{in}C_{out}}.$ (8)
In the above, $\bar{A}^{l}$ is the mean of $A^{l}$, $M$ is a hyperparameter which defines the magnitude of outliers compared to the average activation, and $\mathbf{1}_{\zeta}$ is the indicator function defined as
$\mathbf{1}_{\zeta}=\begin{cases}1&A_{i,j}>M\cdot\bar{A}^{l}\\ 0&\text{otherwise}.\end{cases}$ (9)
For our experimental results, we use the default hyperparameters setting $M=5$ and $\lambda=0.08$.
Our angular distance layer-wise distribution is based on the intuition that parameter allocation to decoder blocks should be based on the propensity of each block to modify the residual stream. In essence, decoder blocks which greatly modify their inputs should be allocated more parameters compared to blocks which only slightly modify their inputs.
Formally, for a decoder with $n$ layers, target global decoder sparsity $S\in[0,1]$, and angular distances $\mathbf{D}\in[D^{1},...,D^{N}]$ where $D^{i}$ is defined by Equation 6 with $n=1$, we allocate the sparsity of each block as
$S^{i}=1-\left((1-S)*\frac{D^{i}}{\sum_{j=1}^{n}\left|D^{j}\right|}\right).$ (10)
(a) Llama-3.2-1B-Instruct
(b) Llama-3.2-3B-Instruct
(c) Llama-3.2-8B-Instruct
(d) Llama-3.2-1B-Instruct
(e) Llama-3.2-3B-Instruct
(f) Llama-3.1-8B-Instruct
Figure 6: Comparison between non-uniform layer-wise sparsity distributions OWL and our proposed angular distance distribution. In Figures 6, 6 and 6, we report the WikiText-V2 Perplexity for Llama-3.2 1B and 3B and Llama-3.1 8B at 50, 60, 70, 80, and 90% sparsity. At high sparsities, the angular distance distribution outperforms OWL, particularly for the 1B model. In Figures 6, 6 and 6, we report the ARC-C accuracy for the same models and sparsities. Despite outperforming OWL in terms of perplexity, we find that OWL achieves slightly higher scores on this benchmark. Further, neither distribution yields statistically significant improvements to acceptance rate in our limited experiments.
Appendix E Benchmarking details
Benchmarking was completed on a 4 $\times$ NVIDIA A6000 GPU node with 88 Intel Xeon E5-2699 v4 CPUs using nm-vLLM (Neural Magic, 2024) version 0.5.0. The throughput of our draft and target models is measured using a single request of 300 input tokens and 256 output tokens, the average token counts from the SpecBench benchmark. We repeat the request 10 times to minimize variance in the measurement. The cost factor, $c$, is calculated as the ratio of the draft and target model tokens per second including both input and output tokens. Using MAL from our SpecBench results, we derive the improvement factor per Equation 3. We report the improvement factor and throughput for dense, layer-pruned, and SD2 unstructured drafters in Figures 1 and 7, respectively. The SD2 1.5B and 3B drafters outperform both dense and layer-pruned models in terms of end-to-end latency improvements.
In Figure 8 we report the improvement factor for 10 concurrent requests following a similar benchmarking approach as per the above. Notably, SD2 still provides the highest improvement factor for the 3B draft model category, but the 1.5B dense drafter slightly outperforms SD2 in this setting.
Figure 7: Throughput (tok/s) vs. sparsity by model and variant. Layer-pruned draft models offer ultra-low latency; however, their reduced MAL results in an overall lower improvement factor compared to SD2 unstructured sparse models. All models are benchmarked using nm-vLLM. Throughput values include both input and output tokens.
Figure 8: Improvement factor of dense, layer-pruned, and SD2 unstructured draft models when benchmarked with 10 concurrent requests. Llama and Qwen models drafting for Llama-3.1-70B-Instruct and Qwen-2.5-72B-Instruct, respectively. Even in this setting, SD2 drafters outperform layer-pruned draft models and dense drafters in the 3B model size category.
Appendix F Detailed SpecBench results
Figure 9: Theoretical improvement factor for SD2 Qwen-2.5 models drafting for Qwen-2.5-72B-Instruct based on MAL and MACs. Similar to our Llama results, SD2 Qwen-3B offers the highest improvement factor compared to dense or layer-pruned models.
In this section, we report the detailed SpecBench (Xia et al., 2024) results for draft models fine-tuned with SD2 (Table 2), one-shot pruned models (Table 3) and SD2 Qwen-2.5 draft models in the UAG setting (Table 4). Figure 9 depicts the theoretical improvement factor for Qwen-2.5 draft models, in which SD2 Qwen-2.5 is found to have the highest improvement factor across all candidate draft models.
Table 2: SpecBench results for SD2 layer-pruned and fine-grained sparse draft models.
Model Variant Sparsity % Overall MT Bench Translation Summarization QA Math Reasoning RAG
Llama-1B Dense 0 $3.92$ $3.99$ $2.97$ $3.43$ $3.55$ $5.65$ $3.66$
Layer-pruned 50 $2.50$ $2.58$ $1.61$ $2.11$ $2.18$ $3.91$ $2.49$
SparseGPT 50 $3.54$ $3.68$ $2.70$ $3.05$ $3.00$ $5.22$ $3.46$
2:4 $3.27$ $3.39$ $2.27$ $2.87$ $2.71$ $5.00$ $3.21$
66 $3.11$ $3.23$ $2.16$ $2.73$ $2.57$ $4.92$ $3.01$
75 $2.98$ $3.12$ $1.97$ $2.50$ $2.49$ $4.84$ $3.01$
Llama-3B Dense 0 $4.54$ $4.58$ $3.59$ $4.12$ $4.26$ $5.88$ $4.27$
Layer-pruned 50 $2.62$ $2.70$ $1.80$ $2.26$ $2.25$ $4.02$ $2.65$
SparseGPT 50 $4.16$ $4.29$ $3.30$ $3.74$ $3.62$ $5.60$ $3.94$
2:4 $3.79$ $3.91$ $2.98$ $3.46$ $3.15$ $5.42$ $3.61$
66 $3.61$ $3.74$ $2.70$ $3.21$ $2.97$ $5.30$ $3.55$
75 $3.31$ $3.47$ $2.27$ $2.85$ $2.71$ $5.14$ $3.24$
Qwen-0.5B Dense 0 $3.36$ $3.44$ $2.99$ $2.81$ $2.58$ $5.70$ $2.85$
Layer-pruned 50 $2.42$ $2.51$ $1.44$ $1.91$ $2.00$ $3.96$ $2.09$
SparseGPT 50 $3.17$ $3.28$ $2.15$ $2.65$ $2.45$ $5.15$ $2.71$
2:4 $2.98$ $3.09$ $1.80$ $2.50$ $2.32$ $4.87$ $2.55$
Qwen-1.5B Dense 0 $3.81$ $3.90$ $3.72$ $3.21$ $2.96$ $6.02$ $3.26$
Layer-pruned 50 $2.62$ $2.73$ $1.59$ $2.08$ $2.16$ $4.16$ $2.31$
SparseGPT 50 $3.61$ $3.75$ $2.83$ $3.07$ $2.82$ $5.35$ $3.13$
2:4 $3.36$ $3.50$ $2.35$ $2.84$ $2.58$ $5.21$ $2.90$
Qwen-3B Dense 0 $4.03$ $4.17$ $3.86$ $3.45$ $3.15$ $6.09$ $3.43$
Layer-pruned 50 $2.74$ $2.87$ $1.58$ $2.16$ $2.24$ $4.38$ $2.37$
SparseGPT 50 $3.87$ $4.04$ $3.47$ $3.36$ $2.97$ $5.46$ $3.41$
2:4 $3.60$ $3.73$ $3.03$ $3.13$ $2.76$ $5.36$ $3.13$
Table 3: SpecBench results for one-shot pruned layer-pruned and fine-grained sparse draft models.
Model Variant Sparsity % Overall MT Bench Translation Summarization QA Math Reasoning RAG
Llama-1B Dense 0 $3.92$ $3.99$ $2.97$ $3.43$ $3.55$ $5.65$ $3.66$
Layer-pruned 50 $1.11$ $1.12$ $1.05$ $1.09$ $1.11$ $1.15$ $1.11$
SparseGPT 50 $2.93$ $3.00$ $2.12$ $2.68$ $2.42$ $4.52$ $2.82$
2:4 $2.27$ $2.26$ $1.58$ $2.23$ $1.91$ $3.29$ $2.33$
66 $1.87$ $1.86$ $1.29$ $1.84$ $1.65$ $2.47$ $1.85$
75 $1.42$ $1.41$ $1.14$ $1.40$ $1.37$ $1.64$ $1.43$
Llama–3B Dense 0 $4.54$ $4.58$ $3.59$ $4.12$ $4.26$ $5.88$ $4.27$
Layer-pruned 50 $1.13$ $1.13$ $1.07$ $1.10$ $1.14$ $1.20$ $1.12$
SparseGPT 50 $3.76$ $3.82$ $2.79$ $3.54$ $3.17$ $5.35$ $3.61$
2:4 $2.87$ $2.85$ $2.04$ $2.92$ $2.32$ $4.35$ $2.93$
66 $2.47$ $2.44$ $1.72$ $2.51$ $2.02$ $3.65$ $2.54$
75 $1.79$ $1.75$ $1.27$ $1.86$ $1.62$ $2.23$ $1.88$
Layer-pruned 50 $1.26$ $1.26$ $1.17$ $1.20$ $1.24$ $1.38$ $1.25$
SparseGPT 50 $2.76$ $2.79$ $2.19$ $2.51$ $2.17$ $4.32$ $2.41$
2:4 $2.21$ $2.21$ $1.63$ $2.18$ $1.84$ $2.90$ $2.04$
Qwen-1.5B Dense 0 $3.81$ $3.90$ $3.72$ $3.21$ $2.96$ $6.02$ $3.26$
Layer-pruned 50 $1.40$ $1.39$ $1.27$ $1.33$ $1.35$ $1.52$ $1.39$
SparseGPT 50 $3.14$ $3.18$ $2.84$ $2.87$ $2.47$ $4.69$ $2.75$
2:4 $2.51$ $2.53$ $1.96$ $2.43$ $2.08$ $3.41$ $2.24$
Qwen-3B Dense 0 $4.03$ $4.17$ $3.86$ $3.45$ $3.15$ $6.09$ $3.43$
Layer-pruned 50 $1.41$ $1.41$ $1.26$ $1.31$ $1.35$ $1.56$ $1.38$
SparseGPT 50 $3.46$ $3.51$ $3.27$ $3.13$ $2.74$ $5.12$ $3.04$
2:4 $2.83$ $2.82$ $2.63$ $2.78$ $2.30$ $3.88$ $2.54$
Table 4: SpecBench results for SD2 layer-pruned and fine-grained sparse draft Qwen models drafting for Llama-3.1-70B-Instruct in the UAG setting.
Model Variant Sparsity % Overall MT Bench Translation Summarization QA Math Reasoning RAG
Qwen-0.5B Dense 0 $2.97$ $3.09$ $2.04$ $2.71$ $2.48$ $4.24$ $2.70$
SparseGPT 50 $3.03$ $3.18$ $2.04$ $2.68$ $2.52$ $4.44$ $2.78$
2:4 $2.87$ $3.01$ $1.77$ $2.54$ $2.38$ $4.35$ $2.68$
Qwen-1.5B Dense 0 $3.34$ $3.47$ $2.34$ $3.09$ $2.82$ $4.59$ $2.97$
SparseGPT 50 $3.42$ $3.58$ $2.28$ $3.08$ $2.92$ $4.70$ $3.04$
2:4 $3.22$ $3.40$ $2.04$ $2.89$ $2.69$ $4.59$ $2.88$
Qwen-3B Dense 0 $3.43$ $3.58$ $2.40$ $3.25$ $2.91$ $4.55$ $2.92$
SparseGPT 50 $3.63$ $3.82$ $2.39$ $3.36$ $3.05$ $4.83$ $3.28$
2:4 $3.44$ $3.60$ $2.28$ $3.17$ $2.87$ $4.73$ $3.13$
Appendix G Detailed OpenLLM Leaderboard V1 results
In Tables 5 and 6, we report the benchmark results for our model’s after fine-tuning and after one-shot pruning, respectively. In Table 7, we report the benchmark suite results for the Qwen draft models used in our UAG experiments.
Table 5: SD2 layer-pruned and fine-grained sparse draft model results on the OpenLLM Leaderboard V1 benchmarks.
Model Variant Sparsity % ARC-C GSM8K Hella -Swag MMLU Wino -grande TruthfulQA MC2 Mean Accuracy
Llama–1B Dense 0 $42.58$ $33.13$ $59.70$ $45.46$ $62.19$ $43.85$ $47.82$
Layer-pruned 50 $25.85$ $0.08$ $34.74$ $25.42$ $52.25$ $42.50$ $30.14$
SparseGPT 50 $37.03$ $21.99$ $52.29$ $33.40$ $58.72$ $45.35$ $41.46$
2:4 $32.34$ $11.98$ $46.23$ $28.36$ $56.99$ $44.04$ $36.66$
66 $30.38$ $7.73$ $40.94$ $26.81$ $55.17$ $43.72$ $34.12$
75 $25.51$ $1.59$ $34.15$ $26.08$ $51.07$ $45.19$ $30.60$
Llama–3B Dense 0 $51.54$ $64.59$ $73.09$ $59.71$ $69.53$ $49.72$ $61.37$
Layer-pruned 50 $27.39$ $0.38$ $36.91$ $25.12$ $50.04$ $44.22$ $30.68$
SparseGPT 50 $43.60$ $50.19$ $65.56$ $52.81$ $65.67$ $49.32$ $54.53$
2:4 $42.75$ $38.21$ $59.93$ $44.50$ $62.59$ $45.77$ $48.96$
66 $40.19$ $29.72$ $54.70$ $39.52$ $60.77$ $48.12$ $45.50$
75 $33.45$ $13.50$ $45.51$ $32.35$ $57.46$ $41.90$ $37.36$
Qwen-0.5B Dense 0 $36.26$ $21.53$ $51.40$ $46.99$ $54.54$ $41.95$ $42.11$
Layer-pruned 50 $25.51$ $0.15$ $31.88$ $25.54$ $51.93$ $46.94$ $30.33$
SparseGPT 50 $32.25$ $5.84$ $44.69$ $38.33$ $54.78$ $43.16$ $36.51$
2:4 $30.03$ $4.09$ $39.98$ $28.24$ $54.22$ $43.50$ $33.35$
Qwen-1.5B Dense 0 $53.92$ $31.31$ $67.70$ $60.35$ $65.59$ $46.61$ $54.25$
Layer-pruned 50 $27.65$ $0.76$ $38.48$ $25.71$ $50.91$ $43.73$ $31.21$
SparseGPT 50 $47.27$ $42.76$ $60.79$ $50.51$ $61.56$ $47.98$ $51.81$
2:4 $39.59$ $26.84$ $54.44$ $40.46$ $58.25$ $43.64$ $43.87$
Qwen-3B Dense 0 $60.24$ $10.54$ $75.18$ $66.37$ $70.40$ $58.76$ $56.92$
Layer-pruned 50 $30.29$ $1.06$ $40.53$ $23.57$ $50.67$ $43.29$ $31.57$
SparseGPT 50 $50.34$ $61.11$ $68.85$ $58.27$ $65.82$ $54.05$ $59.74$
2:4 $46.93$ $44.12$ $62.83$ $49.68$ $65.04$ $50.62$ $53.20$
Table 6: One-shot pruned layer-pruned and fine-grained sparse draft model results on the OpenLLM Leaderboard V1 benchmarks.
Model Variant Sparsity % ARC-C GSM8K Hella -Swag MMLU Wino -grande TruthfulQA MC2 Mean Accuracy
Llama-1B Dense 0 $42.58$ $33.13$ $59.70$ $45.46$ $62.19$ $43.85$ $47.82$
Layer-pruned 50 $26.62$ $0.00$ $27.21$ $22.97$ $48.54$ $47.39$ $28.79$
SparseGPT 50 $34.98$ $4.62$ $48.50$ $31.01$ $59.35$ $40.69$ $36.53$
2:4 $26.71$ $1.06$ $38.15$ $26.15$ $55.72$ $41.85$ $31.61$
66 $24.74$ $0.99$ $32.93$ $26.66$ $52.49$ $41.35$ $29.86$
75 $22.53$ $0.30$ $27.68$ $25.45$ $47.91$ $46.84$ $28.45$
Llama–3B Dense 0 $51.54$ $64.59$ $73.09$ $59.71$ $69.53$ $49.72$ $61.37$
Layer-pruned 50 $25.00$ $0.00$ $28.14$ $24.04$ $49.41$ $49.77$ $29.39$
SparseGPT 50 $42.75$ $31.99$ $62.44$ $49.22$ $63.46$ $47.66$ $49.59$
2:4 $33.62$ $5.46$ $47.78$ $37.48$ $58.17$ $44.72$ $37.87$
66 $28.24$ $1.44$ $39.68$ $30.05$ $57.38$ $43.90$ $33.45$
75 $23.98$ $0.15$ $29.52$ $25.61$ $50.67$ $45.64$ $29.26$
Qwen-0.5B Dense 0 $36.26$ $21.53$ $51.40$ $46.99$ $54.54$ $41.95$ $42.11$
Layer-pruned 50 $21.33$ $0.00$ $27.61$ $22.98$ $50.83$ $49.69$ $28.74$
SparseGPT 50 $29.35$ $3.03$ $42.74$ $33.04$ $53.91$ $42.85$ $34.15$
2:4 $25.09$ $1.44$ $34.61$ $25.92$ $52.57$ $43.50$ $30.52$
Qwen-1.5B Dense 0 $53.92$ $31.31$ $67.70$ $60.35$ $65.59$ $46.61$ $54.25$
Layer-pruned 50 $24.06$ $0.00$ $32.25$ $24.37$ $49.33$ $48.14$ $29.69$
SparseGPT 50 $43.94$ $14.25$ $57.65$ $48.60$ $61.01$ $42.87$ $44.72$
2:4 $31.23$ $2.65$ $43.99$ $35.71$ $56.04$ $40.74$ $35.06$
Qwen-3B Dense 0 $60.24$ $10.54$ $75.18$ $66.37$ $70.40$ $58.76$ $56.92$
Layer-pruned 50 $23.29$ $0.08$ $30.87$ $24.78$ $50.51$ $49.39$ $29.82$
SparseGPT 50 $48.98$ $43.44$ $66.22$ $55.64$ $66.93$ $52.77$ $55.66$
2:4 $39.08$ $8.64$ $53.25$ $43.95$ $60.93$ $42.01$ $41.31$
Table 7: OpenLLM Leaderboard V1 benchmarks for SD2 layer-pruned and fine-grained sparse Qwen-2.5 draft models aligned with a Llama-3.1-70B-Instruct target model for the UAG setting.
Model Variant Sparsity % ARC-C GSM8K Hella -Swag MMLU Wino -grande TruthfulQA MC2 Mean Accuracy
Qwen-0.5B Layer-pruned 50 $26.88$ $0.00$ $32.19$ $25.35$ $50.12$ $45.33$ $29.98$
SparseGPT 50 $32.08$ $6.29$ $44.83$ $38.07$ $55.88$ $44.56$ $36.95$
2:4 $28.84$ $4.55$ $40.50$ $28.00$ $54.70$ $44.34$ $33.49$
Qwen-1.5B Layer-pruned 50 $28.75$ $0.08$ $39.21$ $25.57$ $52.25$ $43.67$ $31.59$
SparseGPT 50 $45.48$ $42.53$ $60.32$ $51.30$ $61.33$ $46.25$ $51.20$
2:4 $38.65$ $28.20$ $53.92$ $40.09$ $58.64$ $43.76$ $43.88$
Qwen-3B Layer-pruned 50 $31.48$ $0.99$ $40.55$ $22.98$ $50.28$ $42.17$ $31.41$
SparseGPT 50 $48.98$ $58.07$ $67.88$ $57.61$ $65.04$ $49.09$ $57.78$
2:4 $44.71$ $45.94$ $61.77$ $48.72$ $63.93$ $48.03$ $52.18$